Three resistors Ra,Rb,& Rc are connected in series to a 120volt source, and the total power taken by the
circuit is 200 watts. Rb = 2Ra, Rc =3Ra. If Ra,Ra,& Rc are connected in parallel, the equivalent resistance
will be?
1
Expert's answer
2021-09-27T18:23:26-0400

Explanations & Calculations


  • Equivalent resistance in the first case is "\small R_e = R_a+R_b+R_c."
  • From the power of the circuit

"\qquad\qquad
\begin{aligned}
\small 200W&=\small \frac{V^2 }{R_e}=\frac{V^2}{R_a+R_b+R_c}\\
\small R_a+R_b+R_c&=\small \frac{120^2}{200}\\
&=\small 72\Omega\cdots(1)
\end{aligned}"


  • When they are connected in parallel the equivalent will be

"\qquad\qquad
\begin{aligned}
\small \frac{1}{R_{e1}}&=\small \frac{1}{R_a}+\frac{1}{R_b}+\frac{1}{R_c}\\
&=\small \frac{1}{R_a}+\frac{1}{2R_a}+\frac{1}{3R_a}\\
\small R_{e1}&=\small \frac{6R_a}{11} \cdots\cdots(2)
\end{aligned}"


  • Now, it is about calculating "\small R_a", which can be done with the help of the equation (1).
  • Give it a try & once you have the value, you can substitute it in equation (2) & get the required answer.